OpenStudy (anonymous):

please help me.. What is the equation of the line that is perpendicular to -x + y = 7 and passes through (-1, -1)?

OpenStudy (igreen):

Perpendicular lines have opposite slopes. -x + y = 7 First, convert to slope-intercept form: -x + y = 7 Add 'x' to both sides: y = x + 7 Now it's in the form of y = mx + b, where 'm' is the slope, so what's the slope here?

OpenStudy (anonymous):

2 i think

OpenStudy (igreen):

No..it would be right before the 'x'..and that is '1'.

OpenStudy (igreen):

The opposite of 1 is -1. So the slope will be -1. Plug this in with the point to get the equation. \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value of the point, \(x_1\) is the x-value of the point, and \(m\) is the slope. So we have: \(y + 1 = -1(x + 1)\) Distribute -1 into the parenthesis: \(y + 1 = -x - 1\) Add 1 to both sides, what's -1 + 1?

OpenStudy (anonymous):

0 right?

OpenStudy (igreen):

Yes, so we have: y = -x

OpenStudy (igreen):

That's our equation.